find the length of the sides of tiangle whose vertices are (1,4),(-4,0),(3,-3).Classify the triangle
Answer by ewatrrr(24785)

 
Hi,
find the length of the sides of tiangle whose vertices are
(1,4),
(-4,0) D = sqrt(4^2 + 5^2) = sqrt(41)
(3,-3 D = sqrt(3^2 + -7^2) = sqrt(58) longest side C
(1,4), D = sqrt(-7^2 + 2^2) = sqrt(53)
41 + 53 = 94 > 58
If A, B and C are the sides of a triangle where C is the longest side, then we can say the following
i) If A^2+B^2 = C^2 is true, then we have a right triangle
ii) If A^2+B^2 > C^2 is the case, then we have an acute triangle.*****
iii) If A^2+B^2 < C^2, then we have an obtuse triangle.